Six square feet of a certain kind of brick wall contains 78 bricks. How many bricks are needed for 150 sq ft?
Answer by Theo(13342)

6 square feet contains 78 bricks.
divide 78 by 6 to get 13 bricks per square foot.
150 square feet would contain 150 * 13 = 1950 bricks.

you can also use a ratio.
78 / 6 = x / 150
solve for x to get x = 78 / 6 * 150 = 1950.